OpenStudy (anonymous):

The fourth root of 32 + the fourth root of 48

OpenStudy (anonymous):

= ( 2 ^(1/4) * 16^(1/4) + (3 ^(1/4) * 16^(1/4) 16^(1/4) = 2 so: = 2*2^(1/4 ) + 2* 3^(1/4) = 2(fourth root 2 + fourth root 3)

OpenStudy (anonymous):

do you need further explanation?

OpenStudy (anonymous):